Using the normal approximation to the binomial to compute the probability of observing 460 or fewer in a sample of 1000 favoring consolidation if we assume that 50% of the entire population favor the change.


1
Expert's answer
2021-07-30T15:16:38-0400

Normal approximation:

"\mu=np=1000*0.5=500,"

"\sigma=\sqrt{np(1-p)}=\sqrt{1000*0.5*0.5}=15.81."

"P(X\le460)=P(X<460.5)=P(Z<\frac{460.5-500}{15.81})=P(Z<-2.50)=0.0062."